OpenStudy (anonymous):

How do you solve : Adam tried to compute the average of his 7 test scores. He mistakenly divided the correct sum of all of his test scores by 6, which yielded 84. What is Adam’s correct average test score?

jimthompson5910 (jim_thompson5910):

Let S = sum of all the scores Since "He mistakenly divided the correct sum of all of his test scores by 6, which yielded 84.", we know that S/6 = 84 Solve for S: S/6 = 84 S = 84*6 S = 504 So the sum of all the scores is 504 To find the correct average, divide the sum S by 7 S/7 = 504/7 = 72 So Adam's correct average test score is 72

OpenStudy (anonymous):

ty

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (precal):

is this the one? I would go with jim's solution. He knows his math